OpenStudy (anonymous):

Find the slope of the line passing through (1, 5) and (0, 2).

OpenStudy (anonymous):

slope = (1-0)/(5-2) = 1/3

OpenStudy (zzr0ck3r):

nono (2-5)/(0-1)

OpenStudy (zzr0ck3r):

upside down:)